# AttnLRP: Attention-Aware Layer-wise Relevance Propagation for Transformers

### v1

@@ 3.3.2 Handling Matrix-Multiplication
The summation already provides a decomposition in the form of equation (1), and we only need to decompose the individual summands $\textbf{A}_{ji}\textbf{V}_{ip}$. Since multiplication is a commutative operation, assigning equal relevance values to both operands is reasonable.
@@
Proposition 3.3 Decomposing matrix multiplication with a sequential application of the uniform rule (14) and the $\varepsilon$-rule (8) yields the following relevance propagation rule:
$R_{ji}^{l-1}(\textbf{A}_{ji})=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R^{l}_{jp}}{2\ \textbf{O}_{jp}+\varepsilon}$ (15)
There is no bias term absorbing relevance. For $\textbf{V}_{ip}$, we sum over the $j$ indices. The proof can be found in Appendix A.3.3. By employing this rule, we are able to maintain strict adherence to the conservation property (3), while also preserving numerical stability.
@@ 3.3.3 Handling Normalization Layers
The work Ali et al., (2022) demonstrates that when linearizing LayerNorm at x, the bias term absorbs most of the relevance equal to $\text{Var}[\textbf{x}]/(\text{Var}[\textbf{x}]+\varepsilon)$, effectively absorbing 99% of the relevance with commonly used values of $\varepsilon=10^{-6}$ and $\text{Var}[\textbf{x}]=1$. Hence, a linearization at x is not meaningful. As a solution, Ali et al., (2022) proposes to skip this operation with the identity rule. In the following, we prove that this heuristic can be derived from the Deep Taylor Decomposition framework.
@@
There is no bias that absorbs relevance. The proof is given in Appendix A.3.4
This rule enforces a strict notion of conservation, while being highly efficient by excluding normalization operations from the computational graph. Experiments in Section 4.1 provide evidence that this simplification is faithful.
@@ 3.4 Understanding Latent Features
Figure 3: There are two approaches for understanding knowledge neurons: (a) Neuron 3948 at the last non-linearity in FFN 17 of the Phi-1.5 model selects a weight row to add to the residual stream. This weight row projected on the vocabulary spans topics about ice, cold places and winter sport. (b) Sentences that maximally activate this neuron contain references about coldness. Attributing the neuron with AttnLRP highlights the most relevant tokens inside the input sentences. Note, although not illustrated here, AttnLRP can also attribute linear layers inside attention modules. Inspired by Voita et al., (2023).
As we iterate through each layer and neuron during the attribution process with AttnLRP, we obtain relevance values for each latent neuron as a by-product, which directly signify their importance for the prediction. Ranking this latent relevance enables us to identify the neurons and layers that are most influential for the reasoning process of the model Achtibat et al., (2023). After identifying a group of neurons, the subsequent step is to reveal the concept that is represented by each neuron.
For Convolutional Neural Networks (CNNs), single units (i.e., neurons or filters) have been shown to correspond to distinct concepts, each fulfilling specific sub-tasks Radford et al., (2017). In order to find the most representative reference samples that explain the neuron’s encoding, researchers rely on Activation Maximization (ActMax) Nguyen et al., (2016), where input samples are sought that give rise to the highest activation value. We follow up on these observations and present the following strategy for understanding latent features:
(1) Collect prompts that lead to the highest activation of a unit. (2) Explain the unit’s activation using AttnLRP, allowing to narrow down the relevant input tokens for the chosen unit.
In this work, we concentrate on knowledge neurons Dai et al., (2022); Voita et al., (2023) that are situated at the last non-linearity in FFN layers $z_{j}=\text{GELU}(W_{1}\textbf{x})$. These neurons possess intriguing properties, as shown in Figure 3: They encode factual knowledge and upon activation, the corresponding row of the second weight matrix $W_{2}$ is added to the residual stream directly influencing the output distribution of the model. By projecting this weight row onto the vocabulary, a distribution of the most probable tokens across the vocabulary is obtained Geva et al., (2022). Applying AttnLRP on ActMax reference samples and projecting the weight row on the vocabulary allow us to understand in which context a neuron activates and how its activation influences the prediction of the next token. In contrast to Ali et al., (2022), AttnLRP also allows analyzing the key and value linear layers inside attention modules.
@@ 4 Experiments
Table 1: Faithfulness scores as area between the least and most relevant order perturbation curves Blücher et al., (2024) on different models and datasets. To assess plausibility, the IoU in the SQuAD v2 dataset is depicted. The higher a score, the better.
Methods ViT-B-16 Llama 2-7b Flan-T5-XL
ImageNet IMDB Wikipedia SQuAD v2 SQuAD v2 IoU
@@ 4.1 Evaluating Explanations (Q1)
A reliable measure of faithfulness of an explanation are input perturbation experiments Samek et al., (2017). This approach iteratively substitutes the most important tokens in the input domain with a baseline value. If the attribution method accurately identified the most important tokens, the model’s confidence in the predicted output should rapidly decrease. The other way around, perturbing the least relevant tokens first, should not affect the model’s prediction and result in a slow decline of the model’s confidence. For more details, see Appendix B.2. Despite its drawbacks, such as potentially introducing out-of-distribution manipulations Chang et al., (2018) and sensitivity towards the chosen baseline value, this approach is widely adopted in the community. Blücher et al., (2024) has addressed this criticism and introduced an enhanced metric by quantifying the area between the least and most relevant order perturbation curves to obtain a robust measure. Hence, we will employ this improved metric to measure faithfulness. Appendix Figure 4 illustrates a typical perturbation curve.
In order to assess plausibility, we utilize the SQuAD v2 Question-Answering (QA) dataset Rajpurkar et al., (2018), which includes a ground truth mask indicating the correct answer within the question. We calculate attributions for accurately answered questions and determine the Intersection over Union (IoU) between the positive attribution values and the ground truth mask. This approach assumes that the model solely relies on the information provided in the ground truth mask, which is not entirely accurate but sufficient for identifying a trend.
@@ 4.1.1 Baselines
We evaluate the faithfulness on two self-attention models, a ViT-B-16 Dosovitskiy et al., (2021) on ImageNet Deng et al., (2009) classification and the Llama 2-7b Touvron et al., (2023) model on IMDB movie review Maas et al., (2011) classification as well as next word prediction of Wikipedia Wikimedia Foundation, (2023). To evaluate cross-attention layers, we use the encoder-decoder model Flan T5-XL Chung et al., (2022) on the SQuAD v2 dataset. We denote our method as AttnLRP and compare it against a broad spectrum of methods including Input×$\times$×Gradient (I×$\times$×G), Integrated Gradients (IG), SmoothGrad (SmoothG), Attention Rollout (AttnRoll), Gradient-weighted Attention Rollout (G×$\times$×AttnRoll) and Conservative Propagation (CP)-LRP. As explained in Appendix A.2.3, we propose to apply the $\gamma$-rule for AttnLRP in the case of ViTs. For better comparison, we also included an enhanced CP-LRP baseline, which also uses the $\gamma$-rule in the ViTs experiment. The LRP variants introduced by Voita et al., (2021); Chefer et al., 2021b are excluded due to numerical instabilities observed in preliminary experiments, see also Appendix A.2.1. Further, we utilize the Grad-CAM adaptation described in Chefer et al., 2021b . Specifically, we weight the last attention map with the gradient. Finally, we expand upon AtMan by incorporating it into encoder-decoder models by suppressing tokens in all self-attention layers within the encoder, while only doing so in cross-attention layers within the decoder. For AtMan, SmoothGrad and Rollout-methods we perform a hyperparameter sweep over a subset of the dataset. More details are in Appendix B.3.
@@ 4.1.2 Discussion
In Table 1, it is visible that AttnLRP consistently outperforms all the state-of-the-art methods in terms of faithfulness, especially in Llama 2. We also observe that gradient-based approaches significantly suffer from noisy attributions, as reflected by the low faithfulness scores and illustrated in example heatmaps in Appendix B.5. CP-LRP with $\varepsilon$ applied on all layers (as proposed in Ali et al., (2022)), also suffers from noisy gradients in ViTs. Applying the $\gamma$-rule for CP-LRP and AttnLRP in ViTs improves the faithfulness substantially. Whereas AtMan does not perform well in unstructured tasks, i.e., next word prediction or classification, it achieves a high score in QA tasks. While G×$\times$×AttnRoll better reflects the model behavior in unstructured tasks compared to AtMan, it is affected by considerable background noise, resulting in a low IoU score in the SQuAD v2 dataset.
@@ 4.3 Understanding & Manipulating Neurons (Q3)
Analogously, for the prompt ‘Children love to eat sugar and’ with the prediction ‘sweets’, the most relevant neuron’s (layer 18, neuron #5687) projection onto the vocabulary signifies a shift in the model’s focus towards the concept of candy, temptation and sweetness in the vocabulary space. We interact with the model by deactivating neuron #3848, and strongly amplifying the activation of neuron #5687 in the forward pass. This manipulation yields the following prediction change:
@@ 5 Conclusion
We have extended the Layer-wise Relevance Propagation framework to non-linear attention, proposing novel rules for the softmax and matrix-multiplication step and providing interpretations in terms of Deep Taylor Decomposition. Our AttnLRP method stands out due to its unique combination of simplicity, faithfulness, and efficiency. We demonstrate its applicability both for LLMs as well as ViTs, utilizing the stabilizing effect of the $\gamma$-rule. In contrast to other backpropagation-based approaches, AttnLRP enables the accurate attribution of neurons in latent space (also within the attention module), thereby introducing novel possibilities for real-time model interaction and interpretation.
@@
To enhance efficiency and reduce memory consumption, custom GPU kernels analogously to FlashAttention Dao et al., (2022) should be implemented for LRP rules. Tuning the $\gamma$-parameter in ViTs to obtain faithful attributions is necessary.
@@ Appendix A Appendix I: Methodological Details
This appendix provides further details on the methods presented in the paper. In particular, we focus on the AttnLRP method and provide implementation details and discuss the stability of the bias term and the noise problem in Vision Transformers. Finally, we provide proofs for the four propositions presented in the main paper.
@@ A.1.3 SmoothGrad
In this work, we set $\mu=0$ and perform a hyperparameter search for $\sigma$ to find the optimal parameter.
@@ A.1.4 Attention Rollout
Self-Attention rollout Abnar and Zuidema, (2020) capitalizes on the intrinsic nature of the attention weights matrix as a representative measure of token importance. It generates a $q\times k$ matrix where each row is normalized to form a probability distribution, representing the importance of each query token to all key tokens. To compute the relevance of hidden layer tokens ($h$) to the original input tokens ($i$), an iterative multiplication of the attention matrices on the left side is sufficient. To account for the residual connection through which the information of the previous tokens flows, an identity matrix I is added:
$\textbf{R}^{h,i}_{k}=(\textbf{I}+\textbf{A}^{h,h})\cdot\textbf{R}^{h,i}_{k-1}$ (23)
where $k=0$ corresponds to the input layer and $R^{h,i}_{-1}$ is initialized with the identity matrix I, $h$ denotes the hidden feature space, and $i$ stands for input dimension.
@@
$\bar{\textbf{A}}=\mathbb{E}_{b}((\nabla\textbf{A}\odot\textbf{A})^{+})$
where $\mathbb{E}_{b}$ denotes the expectation along the head dimension of the attention map. $\bar{\textbf{A}}$ replaces $\textbf{A}^{h,h}$ in (23). For encoder-decoder models, Chefer et al., 2021a present several additional considerations that are not mentioned here for brevity.

### v2

@@ 3.3.2 Handling Matrix-Multiplication
The summation already provides a decomposition in the form of Equation (1), and we only need to decompose the individual summands $\textbf{A}_{ji}\textbf{V}_{ip}$.
@@
Proposition 3.3 Decomposing matrix multiplication with a sequential application of the $\varepsilon$-rule (8) and the uniform rule (14) on the summands yields the following relevance propagation rule for $\textbf{A}_{ji}$:
$R_{ji}^{l-1}=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R^{l}_{jp}}{2\ \textbf{O}_{jp}+\varepsilon}$ (15)
There is no bias term absorbing relevance, whereas $\varepsilon$ absorbs a negligible quantity. For $\textbf{V}_{ip}$, we sum over the $j$ indices. The proof can be found in Appendix A.3.3. By employing this rule, we maintain strict adherence to the conservation property (3), as explained in Appendix A.3.5.
@@ 3.3.3 Handling Normalization Layers
The work Ali et al., (2022) demonstrates that when linearizing LayerNorm at x, the bias term absorbs most of the relevance equal to $\text{Var}[\textbf{x}]/(\text{Var}[\textbf{x}]+\varepsilon)$, effectively absorbing 99% of the relevance with commonly used values of $\varepsilon=10^{-6}$ and $\text{Var}[\textbf{x}]=1$. Hence, a linearization at x is not meaningful. As a solution, Ali et al., (2022) proposes to regard $g(\textbf{x})$ as a constant, which transforms the normalization operation (18) into a (linear) element-wise operation, on which the identity rule (9) can be applied, as discussed in Appendix A.2.2. In the following, we prove that this heuristic can be derived from the Deep Taylor Decomposition framework.
@@
There is no bias that absorbs relevance. The proof is given in Appendix A.3.4. This rule enforces a strict notion of conservation, while being highly efficient by excluding normalization operations from the computational graph. Experiments in Section 4.1 provide evidence that this simplification is faithful.
@@ 3.4 Understanding Latent Features
Figure 3: There are two approaches for understanding knowledge neurons: (a) Neuron 3948 at the last non-linearity in FFN 17 of the Phi-1.5 model selects a weight row to add to the residual stream. This weight row projected on the vocabulary spans topics about ice, cold places and winter sport. (b) Sentences that maximally activate this neuron contain references about coldness. Attributing the neuron with AttnLRP highlights the most relevant tokens inside the input sentences. Inspired by Voita et al., (2023).
As we iterate through each layer during the attribution process with AttnLRP, we obtain relevance values for each latent neuron as a by-product. Ranking this latent relevance enables us to identify neurons and layers that are most influential for the reasoning process of the model Achtibat et al., (2023). The subsequent step is to reveal the concept that is represented by each neuron by finding the most representative reference samples that explain the neuron’s encoding. A common technique is Activation Maximization (ActMax) Nguyen et al., (2016), where input samples are sought that give rise to the highest activation value. We follow up on these observations and present the following strategy for understanding latent features: (1) Collect prompts that lead to the highest activation of a unit. (2) Explain the unit’s activation using AttnLRP, allowing to narrow down the relevant input tokens for the chosen unit.
In this work, we concentrate on knowledge neurons Dai et al., (2022); Voita et al., (2023) that are situated at the last non-linearity in FFN layers $\textbf{z}=\text{GELU}(\textbf{W}_{1}\textbf{x})$. These neurons possess intriguing properties, as shown in Figure 3: They encode factual knowledge and upon activation, the corresponding row of the second weight matrix $\textbf{W}_{2}$ is added to the residual stream directly influencing the output distribution of the model. By projecting this weight row onto the vocabulary, a distribution of the most probable tokens across the vocabulary is obtained Geva et al., (2022). Applying AttnLRP on ActMax reference samples and projecting the weight row on the vocabulary allow us to understand in which context a neuron activates and how its activation influences the prediction of the next token. In contrast to Ali et al., (2022), AttnLRP also allows analyzing the key and value linear layers inside attention modules.
@@ 4 Experiments
Table 1: Faithfulness scores as area between the least and most relevant order perturbation curves Blücher et al., (2024) on different models and datasets. To assess plausibility, the (top-1) accuracy along with the IoU in parentheses are depicted for SQuAD v2. Methods marked with ()($\ast$)( ∗ ) have been proposed here. Additional results for ViT-L-16 and ViT-L-32 are in Appendix Table B.6.
Methods ViT-B-16 LLaMa 2-7b Mixtral 8x7b Flan-T5-XL
ImageNet $\uparrow$ IMDB $\uparrow$ Wikipedia $\uparrow$ SQuAD v2 $\uparrow$ SQuAD v2 $\uparrow$
[…]
@@ 4.1 Evaluating Explanations (Q1)
A reliable measure of faithfulness of an explanation are input perturbation experiments Samek et al., (2017); Hedström et al., (2023). This approach iteratively substitutes the most important tokens in the input domain with a baseline value. If the attribution method accurately identified the most important tokens, the model’s confidence in the predicted output should rapidly decrease. The other way around, perturbing the least relevant tokens first, should not affect the model’s prediction and result in a slow decline of the model’s confidence. For more details, see Appendix B.2. Despite its drawbacks, such as potentially introducing out-of-distribution manipulations Chang et al., (2018) and sensitivity towards the chosen baseline value, this approach is widely adopted in the community. Brocki and Chung, (2023); Blücher et al., (2024) have addressed this criticism and introduced an enhanced metric by quantifying the area between the least and most relevant order perturbation curves to obtain a robust measure. Hence, we will employ this improved metric to measure faithfulness. Appendix Figure B.5 illustrates a typical perturbation curve.
In order to assess plausibility, we utilize the SQuAD v2 Question-Answering (QA) dataset Rajpurkar et al., (2018), which includes a ground truth mask indicating the correct answer within the question. We calculate attributions for accurately answered questions and determine the top-1 accuracy of the most relevant token and the Intersection over Union (IoU) between the positive attribution values and the ground truth mask. This approach assumes that the model solely relies on the information provided in the ground truth mask, which is not entirely accurate but sufficient for identifying a trend.
@@ 4.1.1 Baselines
We evaluate the faithfulness on two self-attention models, a ViT-B-16 Dosovitskiy et al., (2021) on ImageNet Deng et al., (2009) classification and the LLaMa 2-7b Touvron et al., (2023) model on IMDB movie review Maas et al., (2011) classification as well as next word prediction of Wikipedia Wikimedia Foundation, (2023). Additional results for ViT-L-16 and ViT-L-32 are in Appendix Table B.6. To assess plausability, we employ two instruction-finetuned models on the SQuAD v2 dataset: the MoE model Mixtral 8x7b Jiang et al., (2024) and the encoder-decoder model Flan T5-XL Chung et al., (2022). We denote our method as AttnLRP and compare it against a broad spectrum of methods including Input×$\times$×Gradient (I×$\times$×G), Integrated Gradients (IG), SmoothGrad (SmoothG), Attention Rollout (AttnRoll), Gradient-weighted Attention Rollout (G×$\times$×AttnRoll) and Conservative Propagation (CP)-LRP. As explained in Appendix A.2.3, we propose to apply the $\gamma$-rule for AttnLRP in the case of ViTs. For better comparison, we also included an enhanced CP-LRP baseline, which also uses the $\gamma$-rule in the ViTs experiment. The LRP variants introduced by Voita et al., (2021); Chefer et al., 2021b are excluded due to numerical instabilities observed in preliminary experiments, see also Appendix A.2.1. Further, we utilize the Grad-CAM adaptation described in Chefer et al., 2021b . Specifically, we weight the last attention map with the gradient. For a fair comparison, we attribute all methods without the softmax at the classification output, except AtMan which relies on it. KernelSHAP is only evaluated on vision transformers due to prohibitive computational costs on larger LLMs. Finally, we expand upon AtMan by incorporating it into encoder-decoder models by suppressing tokens in all self-attention layers within the encoder, while only doing so in cross-attention layers within the decoder. For AtMan, SmoothGrad and Rollout-methods we perform a hyperparameter sweep over a subset of the dataset. More details about baseline methods and the hyperparameter search are in Appendix A.1 and B.3. We illustrate example heatmaps for SQuAD v2 in Appendix B.7.
@@ 4.1.2 Discussion
In Table 1, we can observe that AttnLRP consistently outperforms all the state-of-the-art methods in terms of faithfulness. In models with a higher number of non-linearities (higher complexity), AttnLRP demonstrates substantially higher accuracy compared to CP-LRP. While the relative improvement to CP-LRP is 3% for Flan-T5-XL, which only utilizes standard attention layers, AttnLRP achieves a remarkable 46% improvement over CP-LRP in terms of top-1 accuracy in Mixtral 8x7b, that incorporates additional expert layers with softmax non-linearities and FFN layers with non-linear weighting. In Appendix B.4, we discuss the architectural differences and conduct an ablation study on different model components to demonstrate this effect. We also observe that gradient-based approaches significantly suffer from noisy attributions, as reflected by the low faithfulness and illustrated in example heatmaps in Appendix B.7. CP-LRP with $\varepsilon$ applied on all layers (as proposed in Ali et al., (2022)), also suffers from noisy gradients in ViTs. Applying instead the $\gamma$-rule for CP-LRP and AttnLRP in ViTs improves the faithfulness substantially. Whereas AtMan and GradCAM do not perform well in unstructured tasks, i.e., next word prediction or classification, they achieve a high score in QA tasks. While G×$\times$×AttnRoll better reflects the model behavior compared to AtMan and GradCAM, it is affected by considerable background noise, resulting in a low IoU score in the SQuAD v2 dataset.
@@ 4.3 Understanding & Manipulating Neurons (Q3)
Analogously, for the prompt ‘Children love to eat sugar and’ with the prediction ‘sweets’, the most relevant neuron’s (layer 18, neuron #5687) projection onto the vocabulary signifies a shift in the model’s focus towards the concept of candy, temptation and sweetness in the vocabulary space. We interact with the model by deactivating neuron #3948, and strongly amplifying the activation of neuron #5687 in the forward pass. This manipulation yields the following prediction change:
@@ 5 Conclusion
We have extended the Layer-wise Relevance Propagation framework to non-linear attention, proposing novel rules for the softmax and matrix-multiplication step and providing interpretations in terms of Deep Taylor Decomposition. Our AttnLRP method stands out due to its unique combination of simplicity, faithfulness, and efficiency. We demonstrate its applicability both for LLMs as well as ViTs, utilizing the denoising effect of the $\gamma$-rule. In contrast to other backpropagation-based approaches, AttnLRP enables the accurate attribution of neurons in latent space (also within the attention module), thereby introducing novel possibilities for real-time model interaction and interpretation.
@@
Adjusting the $\gamma$-parameter in ViTs remains crucial to achieve accurate attributions. To reduce memory consumption, the impact of quantization on attributions and custom GPU kernels for LRP rules should be investigated.
Acknowledgements
We extend our heartfelt gratitude to Leila Arras for her invaluable feedback and to Johanna Vielhaben for improving our faithfulness metric. We also thank Patrick Kahardipraja, Daniel Becking and Maximilian Ernst for their insightful comments.
Impact Statement
This work establishes the foundations that make it possible to systematically analyze and debug transformer-based AI systems, thereby minimizing the occurrence of false or misleading outputs (hallucination) and mitigating biases that may arise from training data or algorithmic processes. Particularly, it opens up the door for future applications of transformer-based AI systems in critical domains such as healthcare and finance, where the ability to explain the model behavior is often a (legal) requirement. The high computational efficiency of our method significantly reduces the energy usage and consequently also the financial overhead and environmental impact associated with the explanation, which will result in a broader adoption of XAI for transformers.
@@ Appendix A Appendix I: Methodological Details
This appendix provides further details on the methods presented in the paper. In particular, we focus on the AttnLRP method, provide implementation details, discuss the stability of the bias term, highlight the difference between other LRP variants, discuss the noise problem in Vision Transformers and illustrate the effects of temperature scaling on attributing the softmax function. Finally, we provide proofs for the four propositions presented in the main paper.
@@ A.1.2 Integrated Gradients
We utilize zennit Anders et al., (2021) and its default settings to compute Integrated Gradients attribution maps i.e. $m=20$.
@@ A.1.3 SmoothGrad
In this work, we set $\mu=0$ and perform a hyperparameter search for $\sigma$ to find the optimal parameter. We utilize zennit Anders et al., (2021) and its default settings to compute SmoothGrad attribution maps i.e. $m=20$.
@@ A.1.4 Attention Rollout
Self-Attention rollout Abnar and Zuidema, (2020) capitalizes on the intrinsic nature of the attention weights matrix $\textbf{A}\in\mathbb{R}^{b\times s_{q}\times s_{k}}$ as a representative measure of token importance. It generates a $s_{q}\times s_{k}$ matrix where each row is normalized to form a probability distribution, representing the importance of each query token to all key tokens. The attention scores along the head dimension are averaged:
$=\mathbb{E}_{b}[\textbf{A}]$ A ]
where $\mathbb{E}_{b}$ denotes the expectation along the head dimension $b$ of the attention map. To compute the relevance of hidden layer tokens ($h$) to the original input tokens ($i$), an iterative multiplication of the attention matrices on the left side is sufficient. Hence, the key dimension represents the inputs and the query dimension the outputs. To account for the residual connection through which the information of the previous tokens flows, an identity matrix I is added:
$\textbf{R}^{h,i}_{k}=(\textbf{I}+\bar{\textbf{A}}^{h,h})\cdot\textbf{R}^{h,i}_{k-1}$ (23)
where $k=1$ corresponds to the input layer and $\textbf{R}^{h,i}_{0}$ is initialized with the identity matrix I, $h$ denotes the hidden feature space, and $i$ stands for input dimension.
@@
$\bar{\textbf{A}}=\mathbb{E}_{b}[(\nabla\textbf{A}\odot\textbf{A})^{+}]$
For encoder-decoder models, Chefer et al., 2021a present several additional considerations that are not mentioned here for brevity.